OpenStudy (anonymous):

An inquisitive physics student and mountain climber climbs a 50.0-m-high cliff that overhangs a calm pool of water. He throws two stones vertically downward, 1.00 s apart, and observes that they cause a single splash. The first stone has an initial speed of 2.00 m/s. (a) How long after release of the first stone do the two stones hit the water? (b) What initial velocity must the second stone have if the two stones are to hit the water simultaneously? (c) What is the speed of each stone at the instant the two stones hit the water?

OpenStudy (anonymous):

Firstly, you need to figure out how long it takes stone 1 to reach the water. Use the formula:\[\Delta y=u_y+\frac{1}{2}at^2\]You know \(\Delta y\) (50.0 m), \(u_y\) (2.00 m/s), and \(a\) (9.81 m/s/s). What's t?

OpenStudy (anonymous):

Sorry, that formula is \[\Delta y=u_yt+\frac{1}{2}at^2\]

OpenStudy (anonymous):

3s

OpenStudy (anonymous):

The second stone was released 1 second after the first stone, and hit the water at the same time, so the total travel time is 2 seconds. Now, using the same formula as before:\[\Delta y=u_{y2}t+\frac{1}{2}at^2\]This time, your unknown is \(u_{y2}\). You have 't', I just told you what that was :)

OpenStudy (anonymous):

Initially that is what I did. However, my answer differs from the books by 0.1 sec. Since i am learning and do not feel confident. i needed to ask

OpenStudy (anonymous):

For a difference of that size, it is possibly a rounding error.

OpenStudy (anonymous):

For example, they might have used 9.8, or 10 as the acceleration due to gravity

OpenStudy (anonymous):

thanks for the assistance